### v2

# Read and Think: An Efficient Step-wise Multimodal Language Model for Document Understanding and Reasoning

Abstract
Understanding the contents of multimodal documents is essential to accurately extract relevant evidence and use it for reasoning. Existing document understanding models tend to generate answers with a single word or phrase directly, ignoring the source document’s evidence and lacking interpretability. In this work, we address the lack of step-wise capabilities through data augmentation and extension. Specifically, We use Multi-modal Large Language Models (MLLMs), which have strong visual understanding and reasoning abilities, as data generators to generate step-wise question-and-answer pairs for document images and use a high-performance LLM as the error detector to filter out noisy data. This step-wise data generation pipeline is implemented using both template-based and few-shot methods. We then use the generated high-quality data to train a humanized document understanding and reasoning model, specifically designed to solve complex questions that require reasoning or multi-hop question answering, dubbed DocAssistant. Experimental results demonstrate the effectiveness and application value of step-wise generation, showing a 5% improvement on InfoVQA with complex layouts and a 7% improvement on ChartQA with complex reasoning, compared to directly generated answers. We hope our work highlights the potential of synthetic data and encourages further exploration of multi-modal document reasoning capabilities.
Instruction
There are various documents in the real world, which differ from images of real-world scenarios. Document images are often filled with extensive text and graph information, requiring the model to have strong capabilities in document layout understanding, text semantic understanding, and numerical reasoning. This presents both a high challenge and high application value. In the document visual question answering (DVQA) task, As shown in Figure 1, existing document understanding models (Xu et al. 2020a, b; Davis et al. 2022) tend to generate answers directly based on questions, making it impossible to determine the source of the answers and their accuracy.
However, we aim for the model to think like a human, first locating the relevant information for questions in documents, and then extracting or inferring an answer based on that. For example, in text-rich documents, this would be the context related to keywords, while in chart documents (Methani et al. 2020; Obeid and Hoque 2020; Han et al. 2023), it would be the relevant values and coordinates. This is a challenging task that requires the model to have both precise identification and strong reasoning abilties.
Recently, a number MLLMs have made breakthroughs in document understanding, such as InternLM-XC2-4KHD (Dong et al. 2024), InternVL (Chen et al. 2024b), Llava-uhd (Xu et al. 2024b), and mPLUG-DocOwl 1.5 (Hu et al. 2024). By improving image resolution and understanding documents from both macro and micro perspectives, these models have greatly enhanced the ability of visual encoders to understand fine-grained information in documents. Most of them have achieved promising performance on document understanding datasets like DocVQA (Mathew, Karatzas, and Jawahar 2021), ChartQA (Masry et al. 2022), InfographicVQA (Mathew et al. 2022), etc. However, they can only handle simple questions in simple-layout documents, such as directly extracting answers from continuous text in the document images. When faced with complex-layout documents, they are likely to make errors in information identification. Additionally, when dealing with complex questions, they tend to ignore the reasoning steps and directly produce answers without providing the source or basis for these answers. Therefore, the existing models have a lot of room for improvement in the questions of documents with complex layouts and documents requiring reasoning.
At present, existing multi-modal document datasets are gradually improving, encompassing a range from the simple layout of scanned documents in DocVQA to the complex layout of poster documents in InfographicVQA, as well as chart documents involving mathematical logic operations in ChartQA. This progression allows for a more comprehensive assessment of the MLLMs’ understanding ability of different types of documents and various questions. However, most of the questions in these labeled datasets are relatively simple, focusing primarily on image information recognition, with fewer questions requiring reasoning. Additionally, the answers do not provide contextual information, making it difficult to intuitively confirm their correctness.
In this paper, We aim to develop an efficient, general-purpose document understanding model that combines document information recognition and reasoning. When answering questions, the model will think like a human: first locating the context information related to the keywords in the question, and then deducing the corresponding answer based on this context information. Since there is no such data at present, we enhance existing document visual question-and-answer data by taking the answer as the supervisory signal and allowing large-scale MLLMs to complete the thinking process from question to answer using our designed templates. Additionally, leveraging the powerful generative capabilities of MLLMs, we can extend the training data through the templates and samples we designed. Since the generated data contains noise, we designed a pipeline based on multi-agent collaborative filtering and rules to filter out these noisy data. Finally, the filtered data is used to train the small-scale vision language model (SLVM) to be efficient and proficient in document understanding and reasoning.
To summarize, our primary contributions include:
We used MLLMs to design a data generator based on templates and few-shots, and a multi-agents-based data filter to augment and extend high-quality, step-wise visual document question-and-answer data.
We trained an efficient SLVM, dubbed DocAssistant, with both question-aware document visual context understanding and reasoning, which has a high practical value.
We achieved state-of-the-art results on complex layout document understanding and reasoning datasets, and provided more extensive experiments to validate the superiority of our model.
Related Work
Visually Rich Document Understanding (VRDU)
The VRDU task is designed to provide a document image and a question, requiring the model to answer the question by understanding the text, images, and layout of the document. Existing document understanding models fall into two categories: OCR-based and end-to-end models that do not rely on external tools. The former includes models like LayoutLMv3 (Huang et al. 2022), UDOP (Tang et al. 2023), and DocFormerv2 (Appalaraju et al. 2024), which use OCR to extract text and corresponding coordinate information and design pre-training tasks based on text, layout, and image mode information to understand documents. The latter includes simpler architectures such as Donut (Kim et al. 2022) and Pix2Struct (Lee et al. 2023), which use only a visual encoder and a text decoder to improve document understanding through pseudo-OCR tasks or pre-training tasks like image masking. With the rise of large language models, the multi-modal architecture model led by LLaVA (Liu et al. 2024) has also made significant breakthroughs. Models (Hu et al. 2024; Dong et al. 2024; Chen et al. 2024a) involving document visual question-answering tasks have achieved good results in simple questions, such as document information extraction with simple layouts, by enhancing the document data of the visual encoder based on this architecture.
However, the above models do not perform well with documents that have a complex layout and questions requiring reasoning. To address this, we introduce a visual chain of thought to train the SVLM by augmenting and extending the existing document visual dataset (Mathew, Karatzas, and Jawahar 2021; Mathew et al. 2022; Masry et al. 2022). This allows them to think like a human when answering questions, thereby improving the accuracy of their responses.
Reasoning step-by-step
For LLMs, chain-of-thought prompting has been found to be a simple and effective method to improve reasoning performance, often used to tackle complex tasks. For example, SymbCoT (Xu et al. 2024a) uses LLMs to implement a system with translation, planning, solving, and verification, following a logical reasoning framework to maximize LLMs’ ability to stimulate the chain of thought. In the multimodal field, recent works such as VisProg (Gupta and Kembhavi 2023) and ViperGPT (Surís, Menon, and Vondrick 2023) utilize LLMs as planners to compose domain-specific models (Lu et al. 2024; Yang et al. 2023; Zeng et al. 2022) to solve complex tasks.
However, in the field of document VQA, there is a lack of data and multimodal models capable of reasoning about document images. Given the powerful capabilities of LLMs in data generation, we employ a comprehensive approach that includes designing templates and exemplars, using tools, etc. This approach involves teaching a smaller model by learning from a larger teacher model to fill gaps in data and modeling for document visual question-and-answer tasks.
Method
Given a document image and a question, the goal is to enable the model to think and answer step by step like a human. First, design a data generator and a data filter to construct high-quality data. These data are then used to train an efficient context-interpretable multi-modal document question-answering model.
In this section, We first describe the dataset construction process. Next, we explain how to filter out noisy data to retain only high-quality data. Finally, we describe our model setup, using these data to train a dedicated multi-modal document question-answering model with contextual reasoning.
Data Generator and Checker
The overview of our MLLM-based data generator and checker is shown in Figure 2. We leverage InternVL-1.5 (Chen et al. 2024a) as the MLLM for data generation and the LLaMA3-70B for data check. Data generation includes creating data based on the existing training set and generating triplets (question, rationale, answer) through our designed templates, which can be found in Appendix A. During the generation process based on the existing training set, we generate a related rationale using the template , image , question , and answer as inputs:
During the triplet generation process, we design a question generation template. In particular, we generate prompt templates for abstracted or abstract documents and use few-shot learning for chart documents:
During data inspection, we use tools to extract text information from documents, specifically using OCR tools for general documents and DEPLOT (Liu et al. 2022) to convert charts into text-based tables. By designing error detection templates, the LLM can determine whether the corresponding logical step is correct based on the problem and the gold text of the image. This includes (1) the extraction of information step from the document itself and (2) the reasoning steps based on the extracted information. If any judgment errors are made, it will return False and delete the piece of data.
Model Archecture
The architecture of our trained model is shown in Figure 3. We leverage Mini-InternVL-Chat-2B as the backbone for Document VQA. This model uses a pre-trained ViT (Dosovitskiy et al. 2020) to extract image features, a trainable MLP module to map extracted image features to the text embedding space, and InternLM2-Chat-1.8B (Zheng Cai and Chen 2024) as the language model to understand user instruction.
Relying on pre-trained visual encoders alone to understand text-heavy images is insufficient. Therefore, external tools can be used to extract text information from the document as a supplement, input it into the language model, and ensure that the context information understood by the model is accurate.
Extractive/Abstractive Answer Generation
For the query document, the answer comes directly from certain parts of the document, and processing this answer generally requires two steps. First, it is necessary to locate the context information related to the keywords of the question, which may be one or more fragments, and then deduce the final answer based on this context information. That is, based on the keywords of the question (Q), the relevant contextual information fragments (C) are located in the document (D). Assuming we use the function ExtractContext to extract this context information, it can be expressed as:
where (C) represents the collection of contextual information fragments relevant to the question (Q). Next, based on the extracted context information (C), the final answer (A) is deduced. Suppose we use the function InferAnswer to infer the answer, then it can be expressed as:
Reasoning Answer Generation
Currently, existing reasoning document images mainly focus on numerical calculations, with the human dataset in ChartQA being a typical example. The answers are not directly present in the documents, they need to be deduced based on the data in the documents. The entire process includes multi-step calculations, where complex nested operations are decomposed to obtain more accurate answers. Since the number of steps is uncertain, we design few-shot templates during the training data generation process, enabling the model to respond flexibly to each question and generate corresponding reasoning steps. The whole process is as follows:
Given a document containing data (D), the question (Q), and the answer (A).
Flexibly respond to each question: design a few-shot template so that the model can flexibly generate reasoning steps for each question. Each step can be represented as:
For each question (Q), the model generates corresponding reasoning steps until the final answer A is obtained.
Training with rationales
Due to the lack of data, especially since ChartQA’s human training set has only 7,398 question pairs, we expanded the data in the original training set and developed a template for question generation to produce different types of data. We also filter noise data using our designed checker.
Experiment
Experimental Setup
Dataset
We run experiments on three document VQA datasets: DocVQA (Mathew, Karatzas, and Jawahar 2021), InfographicVQA (Mathew et al. 2022), and ChartQA (Masry et al. 2022). DocVQA images come from scanned documents, including letters, tables, articles, etc. Most of the questions are relatively simple and involve text extraction directly. InfographicVQA images are taken from posters, where the layout is complex. Most of the questions involve text extraction, and some require simple reasoning. The charts in the ChartQA dataset include bar charts, line charts, and pie charts, all of which are synthetic data or human-marked data. Most of the synthetic data are simple chart information identification, while the human-marked data contain complex mathematical logic operations.
Generated Data
Table 1 shows the generated QA statistics. The sum includes data based on the original training set and data generated from the question generation template extension. The template can be found in the Appendix (Table 3,4,5). For DocVQA, 3 new questions were generated for each image, while InfographicVQA had 4 new questions per image. Given the small amount of original data in ChartQA Human, Since there is less raw data in ChartQA Human, we collected the related Chart images of Chart-to-text (Kantharaj et al. 2022) and the corresponding gold table to generate more chart reasoning data similar to human. Filtered indicates the data filtered through the LLM and rules.
In addition, we generated questions based on the types of questions in each dataset, such as Count, Spatial, and Reasoning data for DocVQA. For more details, see the Appendix (Table 6 and Table 8).
| Dataset | Images | Generated | Filtered |
| DocVQA | 10194 | 39459+30582 | 58324 |
| InfoVQA | 4406 | 23945+17624 | 36832 |
| ChartQA | 18317+44096 | 7398+80831 | 67649 |
Compared Methods
We compare existing document understanding models across various categories: (1) plain text models represented by T5, (2) the LayoutLM series represented by LayoutLMv3, and DocFormerv2, which has the best performance among the T+L+V mode models, (3) the first OCR-free model Donut for understanding documents with image encoders, and Pix2Struct for performing best in end-to-end small-scale document understanding models, and (4) multimodal document understanding models combined with LLMs.
Implementation Details
During the training process, we set 2 epochs with a batch size of 8 and a learning rate of 4e-5. Specifically, dynamically resizing the image, adjusting the resolution to 448x448, and setting the maximum patch size to 12 are crucial for understanding document-type images. All experiments were conducted on a 2B model, with training focused solely on the projection layer and the language model incorporating LoRA.
Main Results
| Method | Modality | Param | DocVQA | InfographicVQA | ChartQA |
| ANLS | ANLS | Relax Accuracy | |||
| T5 (Raffel et al. 2020) | T | 223M | 70.4 | 36.7 | 59.8 |
| LayoutLMv3 (Huang et al. 2022) | T+V+L | 368M | 83.37 | 45.1 | - |
| DocFormerv2 (Appalaraju et al. 2024) | T+V+L | 750M | 87.84 | 48.8 | - |
| Donut (Kim et al. 2022) | V | 201M | 67.5 | 11.5 | 41.8 |
| Pix2Struct (Lee et al. 2023) | V | 1.3B | 76.6 | 40.0 | 58.6 |
| UReader (Ye et al. 2023) | V | 7B | 65.4 | 42.2 | 59.3 |
| Qwen-VL (Bai et al. 2023) | V | 7B | 65.1 | 29.9 | 65.7 |
| mPLUG-DocOwl (Hu et al. 2024) | V | 7B | 62.2 | 38.2 | 57.4 |
| SMoLA-PaLI-3 (Wu et al. 2024) | V | 5B | 84.5 | 52.4 | 68.9 |
[…]
| With gold text pipeline | |||||
| SMoLA-PaLI-3 (Wu et al. 2024) | T+V | 5B | 87.4 | 57.3 | 68.9 |
| DocAssistant(Ours) | T+V | 2B | 88.12 | 65.72 | 78.28 |
Comparison with Existing Models
Existing models only provide a single word or phrase as an answer for the DVQA task. On one hand, it is impossible to determine its source or verify its correctness. On the other hand, for complex questions requiring reasoning, ignoring intermediate reasoning steps is more likely to lead to errors. Through our reconstruction of the dataset and efficient training on the 2B multimodal model, we have achieved impressive results. Using only a small number of parameters, it outperforms many larger document understanding models and demonstrates greater reasoning power than other smaller models.
In terms of details, DocAssistant’s performance on DocVQA is inferior to DocFormerv2, indicating that understanding multimodal information of documents is more accurate than understanding only picture information. However, the existing model performs poorly for documents with complex layouts like InfoVQA and documents with complex questions requiring reasoning. In particular, small-scale models have little or no document reasoning ability. We have addressed this gap while surpassing larger models (¡=7B) in document understanding and reasoning.
Comparison of Different Strategies
| Strategy | DocVQA | InfoVQA | ChartQA | |
| aug. | human | |||
| single-word | 81.02 | 49.38 | 68.96 | 51.04 |
| zero-shot | 79.58 | 53.43 | 72.8 | 52.88 |
| few-shot | 78.13 | 51.09 | 77.68 | 54.96 |
| finetune | 81.83 | 54.64 | 81.28 | 62.32 |
In the experiment of strategy in Table 3, the zero-shot setting involved changing the original instruction to ”Answer the question step by step”. According to the experimental results, the model performs better on complex layout documents and complex reasoning questions when answering questions step by step. For extracted answers to simple documents like DocVQA, the results are similar. For the few-shot setup, we used 3-shot in the experiment and found that the instruction compliance of the 2B model was poor. Its answers did not follow the example format at all, especially in DocVQA and InfoVQA, which did not require reasoning and had a rather casual answer style. Although ChartQA did not follow the example format, it still answered step by step, leading to improved performance, while the other two datasets with extracted answers performed poorly.
Single-word answers can more directly fit the answer, but they lack contextual information, are less interpretable, and cannot be judged for correctness. Zero-shot and few-shot are unstable and heavily influenced by instruction settings. For documents with extractive or abstractive answers, most answers still do not provide source information. These are the original intention behind our approach to building step-by-step datasets and training the model. The results of fine-tuning show that the model’s reasoning ability has greatly improved. Not only is the answer accuracy higher but the source information of the answer can also be provided, helping users confirm the reliability of the answer.
Ablations and Analysis
| DocVQA | InfoVQA | ChartQA | |
| before_extend | 83.15 | 59.64 | 72.32 |
| after_extend | 84.48 | 60.87 | 74.25 |
| after_check | 85.58 | 62.45 | 77.44 |
The Effective of Extend Data
The modification of the original training set data, combined with our extended data, is sufficient for the model to learn the required response style. Experimental results show a performance gain after extension. This also provides a data generation tool for other smaller models, enabling them to further improve performance by generating more data.
The Effective of Checker
We fine-tuned the model using both pre-filtered and filtered data and the results showed improved performance on all three datasets, especially InfoVQA. This indicates that noisy data significantly interfered with documents with complex layouts and proves the effectiveness of our designed checker.
The Effective of Extend Data
The modifications based on the original training set data, combined with our extended data, are sufficient for the model to learn the required response style. As shown in the experimental results in Table 4, the performance improves. This also provides a data generation tool for other smaller models, enabling further performance improvements by generating more data.
Qualitative Analysis
Figure 4 shows a comparison between cases generated by DocAssistant and those generated by other models (T, T+L+V, V). The answers provided by other models in different documents are all single words or phrases, which tend to be incorrect when faced with questions such as chart reasoning and document information extraction with complex layouts, and they do not contain contextual information. The first two examples are questions and answers about complex layout documents, where DocAssistant provides detailed context. The third example is a chart reasoning question, where DocAssistant also gives specific reasoning steps, providing stronger interpretation and higher accuracy and allowing users to understand how the model reached the answer and easily identify any mistakes. Most importantly, model performance can be further improved through extensive training with step-wise data.
However, due to the model’s inherent limitations, DocAssistant still makes mistakes in reasoning on complex questions. Related error case studies can be found in the Appendix Figure 1 and Figure 2.
Analysis of Different Question Types
Figure 5 shows the proportion of question types in each of the three datasets. The Text_extractive type has the largest proportion in all three datasets, especially in DocVQA. InfoVQA and ChartQA have a higher proportion of Count and Reasoning data, respectively, where DocAssistant shows more advantages.”
Table 5 shows how five types of questions in three different types of datasets are represented in different strategies. The experimental results show that the SFT strategy (DocAssistant) has the best performance for all count and reasoning questions in the three datasets. In addition to DocVQA (simple layout, simple question), zero-shot and few-shot performed better than the straightforward setup on InfoVQA (complex layout) and ChartQA (complex question), indicating that the chain-of-thought is effective in solving questions of such documents. It is worth noting that due to the small amount of color and spatial data, there is a large variance in different method settings. For example, ChartQA only has 43 Color-type questions and 5 Spatial-type questions. Interestingly, both few-shot and step-wise training can reduce the comprehension of color information, The possible reason is that the visual encoder’s recognition performance of color information is unstable.
| Dataset | Method | Color | Text | Spatial | Count | Reasoning |
| DocVQA | SF | - | 81.65 | 85.46 | 80.04 | 51.69 |
| ZS | - | 79.77 | 83.48 | 81.0 | 61.48 | |
| FS | - | 78.33 | 79.24 | 81.43 | 55.28 | |
| SFT | - | 82.03 | 84.53 | 81.24 | 65.72 | |
| InfoVQA | SF | 44.83 | 50.32 | 28.57 | 49.13 | 43.37 |
| ZS | 55.17 | 48.65 | 30.0 | 58.85 | 74.12 | |
| FS | 51.72 | 47.82 | 25.0 | 54.04 | 70.24 | |
| SW | 58.62 | 48.76 | 30.0 | 59.84 | 77.01 | |
| ChartQA | SF | 67.44 | 76.05 | 75.0 | 52.50 | 50.74 |
| ZS | 65.12 | 77.90 | 87.50 | 54.37 | 55.09 | |
| FS | 32.56 | 79.17 | 12.5 | 60.45 | 51.29 | |
| SFT | 39.53 | 78.80 | 1.0 | 71.65 | 70.61 |
Conclusion
This paper introduces a new model named DocAssistant, designed to perform document understanding and reasoning. Specifically, we first transform the data of the existing document training set to include the intermediate analysis process in answering questions and then expand the training data by adding missing types such as Spatial, Count, and Reasoning, a filter is designed during this period for generated data, resulting in high-quality data that can then be trained using an efficient 2B model. The results surpass existing larger-scale models and are state-of-the-art across three datasets. Additionally, we conducted extensive experiments, comparing different strategies of the model, performance before and after data expansion and filtering, and various types of questions to validate our work.
However, the current general document visual question answering data for reasoning is still scarce, and SVLMs also have shortcomings in understanding document elements such as color and spatial information. In the future, we aim to address these deficiencies and hope our work will contribute to advancements in the field.